### v1

@@ 3.3. Stage II: Local Enhancement Model
(4) $X=\text{Concat}(C_{g}\otimes C_{m},\ C_{l}),\quad X\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{4}},\vspace{-0.1cm}$
@@
(6) $x_{t}=\sqrt{\bar{\alpha}_{t}}\cdot x_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I}),\vspace{-0.1cm}$
@@
(7) $Z=\text{Concat}(x_{t},\ C_{M},\ X),\quad Z\in\mathbb{R}^{9\times\frac{H}{8}\times\frac{W}{4}}.\vspace{-0.1cm}$
@@
3.4. Training and Inference
Training. The training process is divided into two stages, each targeting a specific set of objectives with separate optimization strategies. Both stages adopt mean squared error (MSE) loss to supervise the denoising process.
Stage I. The global appearance model $\epsilon_{\theta_{g}}$ is trained to synthesize garments that align with the target silhouette and color under textual guidance. To preserve the generative capacity of the pretrained denoising UNet, we freeze all parameters except those of the silhouette UNet and the cross-attention projections in the mixed attention module. Given silhouette features $C_{s}$, text embeddings $C_{t}$, and color features $C_{c}$, we adopt a decoupled training strategy with $L_{\text{silhouette}}$ and $L_{\text{color}}$ losses:
(8) $L_{\text{silhouette}}$ $=\mathbb{E}_{x_{0},\epsilon,C_{t},C_{s},t}\left\|\epsilon-\epsilon_{\theta_{g}}(x_{t},C_{t},C_{s},t)\right\|^{2},$
@@ 3.4. Training and Inference
Stage II. The local enhancement model $\epsilon_{\theta_{l}}$ refines the coarse latent by injecting logos at user-defined locations. To reduce overhead, we fine-tune only the self-attention layers of the logo UNet. Given logo feature $C_{l}$, spatial mask $C_{m}$, and garment latent $C_{g}$, the training objective $L_{\text{logo}}$ is:
(9) $L_{\text{logo}}=\mathbb{E}_{x_{0},\epsilon,C_{l},C_{m},C_{g},t}\left\|\epsilon-\epsilon_{\theta_{l}}(x_{t},C_{l},C_{m},C_{g},t)\right\|^{2},\vspace{-0.1cm}$
@@
Inference. IMAGGarment-1 supports end-to-end inference through a two-stage pipeline operating in a shared latent space. The global appearance model first generates a latent of coarse garment image conditioned on the input text prompt, silhouette, color, and mask. This process is guided by classifier-free guidance (CFG) (Ho and Salimans, 2022):
(10) $\breve{\epsilon}_{\theta_{g}}(x_{t},C_{t},C_{s},C_{c},t)$ $=\ w\cdot\epsilon_{\theta_{g}}(x_{t},C_{t},C_{s},C_{c},t)$
$\quad+(1-w)\cdot\epsilon_{\theta_{g}}(x_{t},t),$
@@
(11) $\breve{\epsilon}_{\theta_{l}}(x_{t},C_{l},C_{m},C_{g},t)$ $=\ w\cdot\epsilon_{\theta_{l}}(x_{t},C_{l},C_{m},C_{g},t)$
$\quad+(1-w)\cdot\epsilon_{\theta_{l}}(x_{t},C_{m},C_{g},t).$
4. Experiments
4.1. Dataset and Metrics
Figure 4. (a) Data construction pipeline for GarmentBench. (b) Example samples with multimodal annotations: silhouette, logo, text, logo location, and color.
Dataset Construction. As shown in Figure 4 (a), we construct and release GarmentBench, a large-scale dataset for fine-grained garment generation, containing multi-modal design conditions such as text, sketches, colors, logos, and location masks. It serves as a controllable and extensible benchmark for advancing personalized fashion generation. The construction process is as follows:
(1) Image Collection and Preprocessing. We collect over 189K high-quality garment images from the internet, covering a wide range of categories such as tops, bottoms, and dresses. To eliminate background distractions and focus on the garment region, we apply YOLOv8 (Jocher et al., 2023) for clothing detection and perform tight cropping to obtain clean garment-centric images for further processing.
(2) Text, Sketch, and Color Extraction. For each image, we automatically generate three auxiliary conditions to simulate real-world design guidance: textual descriptions generated by the multi-modal LLM Qwen-VL-Chat (Bai et al., 2023), covering key attributes such as color, silhouette, and style; structural sketches obtained using Informative-Drawings (Chan et al., 2022), providing shape and layout priors; and color palettes extracted from single-color garments identified via ResNet50 (He et al., 2015) and clustered using K-means (MacQueen, 1967).
@@ 4.1. Dataset and Metrics
As shown in Figure 4 (b), we present representative samples from the GarmentBench dataset, which include fine-grained garment images paired with multi-modal conditions such as textual descriptions, structural silhouettes, color references, logos, and spatial location masks. Additionally, we randomly sample images from the Fashion-ControlNet-Dataset-V3111https://huggingface.co/datasets/Abrumu/Fashion_controlnet_dataset_V3 and apply the same preprocessing pipeline as GarmentBench to construct a test set with 1,267 image-condition pairs for evaluation and comparative analysis.
@@
Evaluation Metrics. We adopt five metrics to comprehensively evaluate visual quality, conditional consistency, and fine-grained controllability. Fréchet inception distance (FID) (Heusel et al., 2017) measures the distribution similarity between generated and real images, reflecting overall realism. Learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) reflects human-perceived visual similarity, effectively capturing structural and textural consistency. Color structure similarity (CSS) (Zeng et al., 2014) assesses the consistency of color distribution, measuring color controllability. Logo location accuracy (LLA) (Fujitake, 2024) quantifies the spatial deviation between generated and target logo positions, reflecting spatial precision. Lastly, CLIPScore (Hessel et al., 2021) measures semantic alignment between the generated image and the input text prompt. These metrics comprehensively assess quality, controllability, and semantic alignment in fine-grained garment generation. See supplementary material for computational details.
4.2. Implementation Details.
@@
In our experiments, both the silhouette UNet and the denoising UNet are initialized with the pretrained Stable Diffusion v1.5 model [^2] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5. The local enhancement model is based on the inpainting variant of Stable Diffusion v1.5333https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-inpainting, with only the self-attention layers being fine-tuned to reduce computational cost. We adopt OpenCLIP ViT-H/14444https://github.com/mlfoundations/open_clip as the CLIP image encoder. All input images are resized to $512\times 640$ resolution. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a constant learning rate of $1\times 10^{-5}$. The global appearance model and the local enhancement model are trained for 150K and 50K steps, respectively, using a batch size of 20. During inference, we adopt the DDIM sampler (Song et al., 2022) with 50 sampling steps. Unless otherwise specified, the silhouette weight $\alpha$ and color weight $\beta$ in Eq.1 and Eq.3 are set to 0.6 and 1.0. The classifier-free guidance (CFG) scale $w$ in Eq.10 and Eq.11 is set to a default value of 7.0.
Figure 5. Qualitative results on seen and unseen GarmentBench samples. The seen set uses original test pairs, while the unseen set involves randomly mixed conditions. IMAGGarment-1 delivers the most consistent outputs, achieving accurate silhouette, color, and logo control across both settings.
4.3. Baseline Comparisons
Due to the absence of prior work tailored to fine-grained garment generation with multi-condition control, we compare our method against four representative baselines: BLIP-Diffusion (Li et al., 2023), AnyDoor (Chen et al., 2024b), ControlNet (Zhang et al., 2023a), and IP-Adapter (Ye et al., 2023). For subject-driven generation methods, BLIP-Diffusion (Li et al., 2023) leverages a learnable Q-Former to align textual and visual embeddings in the latent space, initially designed for subject-preserving generation from text-image pairs. AnyDoor (Chen et al., 2024b) combines identity and detail encoders to reconstruct personalized content, which we adapt to conditions of garment appearance and logo inputs. For plugin-based baselines, we extend ControlNet (Zhang et al., 2023a) and IP-Adapter (Ye et al., 2023) by duplicating and modifying their conditional branches to support multi-conditional inputs, such as silhouette, color, and logo. The adapted versions are referred to as ControlNet-Garment and IP-Adapter-Garment. To ensure task relevance, all methods are fine-tuned on our GarmentBench dataset with support for logo-specific conditioning. All methods are trained and evaluated under identical training protocols, input resolutions, and hardware setups. The corresponding quantitative and qualitative results are presented in Table 1 and Figure 5, respectively, with detailed analysis provided below.
Quantitative Results. As shown in Table 1, IMAGGarment-1 achieves the best performance across all five metrics on the GarmentBench dataset, demonstrating its superiority in controllable fine-grained garment generation. Compared to subject-driven methods (BLIP-Diffusion (Li et al., 2023), AnyDoor (Chen et al., 2024b)), which rely on global features for personalized reconstruction, IMAGGarment-1 shows substantial improvements in FID, LPIPS, and CSS. These gains highlight the effectiveness of our mixed attention and color adapter modules in achieving coherent multi-condition fusion, resulting in more realistic, perceptually consistent, and color-faithful outputs. In contrast to plugin-based approaches (ControlNet-Garment (Zhang et al., 2023a), IP-Adapter-Garment (Ye et al., 2023)) that simply stack independent conditional branches, IMAGGarment-1 yields significantly higher LLA and CLIPScore, reflecting more precise logo placement and stronger semantic-text alignment. Our proposed A3 module drives these improvements, which adaptively injects spatial priors and logo features into the latent space for accurate local control. Overall, these results indicate that global-only conditioning or naive plugin stacking is insufficient for fine-grained control. By contrast, IMAGGarment-1 provides an effective solution for multi-conditional garment synthesis, enabling precise coordination of global structure and local detail.
@@ 4.3. Baseline Comparisons
Table 1. Quantitative comparisons on GarmentBench. Ours achieves the top results across all metrics, with best in bold.
Method FID $\downarrow$ LPIPS $\downarrow$ CSS $\downarrow$ LLA $\uparrow$ CLIPScore $\uparrow$
BLIP-Diffusion* (Li et al., 2023) 92.5187 0.6766 104.8677 0.1677 0.3144
AnyDoor* (Chen et al., 2024b) 60.0643 0.1728 46.6839 0.7487 0.2836
ControlNet-Garment* (Zhang et al., 2023a) 31.7169 0.3978 82.0181 0.5304 0.3449
IP-Adapter-Garment* (Ye et al., 2023) 24.5061 0.3416 76.4631 0.6609 0.3295
IMAGGarment-1 (Ours) 11.7038 0.0904 33.1443 0.7336 0.3460
@@
Qualitative Results. Figure 5 presents qualitative comparisons on both seen and unseen garments. Notably, the seen test set refers to the designated test split of our GarmentBench dataset. To assess generalization, we construct an unseen test set by randomly recombining input conditions (e.g., silhouette, color, logo) to simulate real-world fashion design scenarios. On seen garments, subject-driven methods (BLIP-Diffusion (Li et al., 2023), AnyDoor (Chen et al., 2024b)) reconstruct global appearance but lack spatial control. BLIP-Diffusion retains logo identity yet fails at precise placement due to text-only conditioning, while AnyDoor introduces logo distortions and stylistic artifacts. Plugin-based baselines (ControlNet-Garment (Zhang et al., 2023a), IP-Adapter-Garment (Ye et al., 2023)) treat conditions independently, resulting in poor coordination. ControlNet-Garment suffers from cross-condition interference, and IP-Adapter-Garment often misplaces logos despite preserving structure. In contrast, IMAGGarment-1 achieves accurate control over silhouette, color, and logo placement. On unseen garments, all baselines degrade notably. Subject-driven methods fail to generalize to novel layouts, AnyDoor distorts appearance, and BLIP-Diffusion struggles with logo positioning. Plugin-based methods also falter: ControlNet-Garment produces mismatched outputs, and IP-Adapter-Garment cannot interpret unseen spatial semantics. IMAGGarment-1 remains robust, maintaining alignment across all conditions. This generalization stems from our $A^{3}$ module, which effectively integrates spatial and visual cues in the latent space. These results validate the controllability and flexibility of our method in both seen and unseen settings.
Table 2. Quantitative ablation results on GarmentBench.
Method FID $\downarrow$ LPIPS $\downarrow$ CSS $\downarrow$ LLA $\uparrow$ CLIPScore $\uparrow$
B0 130.1314 0.6495 105.1591 0.1719 0.3476
B1 44.2250 0.1633 34.4135 0.2997 0.3145
B2 23.9822 0.3336 97.245 0.6099 0.3454
B3 20.0446 0.1488 62.2521 0.6701 0.3413
B4 37.8463 0.3881 101.8441 0.5780 0.3359
Full 11.7038 0.0904 33.1443 0.7336 0.3460
4.4. Ablations and Model Analysis
Ablation of Architecture Design. To validate the effectiveness of each component in our framework, we design a series of ablation variants within the IMAGGarment-1 architecture: B0 uses the vanilla Stable Diffusion v1.5 without any of our proposed modules, serving as the baseline. B1 removes the local enhancement model (Stage II), evaluating the impact of omitting logo injection and spatial control. B2 removes the global appearance model (Stage I), assessing the model’s performance without structured silhouette and color conditioning. B3 removes the color adapter from the global appearance model, isolating the role of color guidance in generation. B4 replaces our mixed attention with vanilla self-attention in the denoising UNet, testing the importance of spatial fusion with silhouette features. Full represents the complete IMAGGarment-1 framework with all proposed modules integrated.
Table 2 presents the quantitative impact of each component in our proposed IMAGGarment-1. In B1, which removes the local enhancement stage, the model struggles to place logos precisely, leading to degraded LLA and CLIPScore. Although the overall garment structure is preserved, the lack of spatial control prevents accurate logo integration. In B2, without the global appearance stage, the model fails to maintain silhouette and color consistency, resulting in significantly worse FID, LPIPS, and CSS. This demonstrates that local injection alone is insufficient to handle global garment layouts. B3 disables the color adapter, causing notable drops in CSS, highlighting its role in faithful color transfer and control. B4 replaces our mixed attention with standard self-attention, which weakens the fusion of silhouette guidance and causes drops in both LPIPS and FID, indicating reduced realism and structural coherence. The full IMAGGarment-1 achieves the best performance across all metrics, validating the complementary design of each module’s effectiveness in handling multi-condition garment generation. Further, Figure 6 shows qualitative comparisons. B1 fails to align logos spatially, while B2 produces distorted garments lacking color and silhouette guidance. Despite maintaining logo placement, B3 leads to color mismatch, and B4 generates less coherent garment layouts. In contrast, the full model successfully synthesizes garments with accurate silhouettes, precise logo placement, and faithful color reproduction, demonstrating the benefits of our dual-stage design, color adapter, and mixed attention fusion.
Figure 6. Qualitative ablation results on GarmentBench.
@@ 4.4. Ablations and Model Analysis
Figure 7. Controllability visualization. Each block varies one input condition while keeping others fixed. Top: Silhouette changes lead to consistent structural adaptation. Middle: Color palette variation results in accurate color transfer. Bottom: Logo mask adjustment yields precise spatial placement.
Controllability Analysis. To evaluate the controllability of the proposed IMAGGarment-1, we conduct conditional manipulation experiments by varying one condition at a time, silhouette, color palette, or logo position, while keeping the others fixed. As shown in Figure 7, each two-row block highlights the model’s response to a specific condition. When changing the silhouette (top block), the generated garments conform precisely to the target shapes, demonstrating the capability of the mixed attention module to preserve structural alignment. With color palette variation (middle block), the model consistently reproduces the desired color distributions, verifying the effectiveness of the color adapter in enabling color-faithful generation. Finally, adjusting the logo position (bottom block) yields accurate spatial relocation, highlighting the $A^{3}$ module’s ability to inject spatial priors for precise local control. These results demonstrate that IMAGGarment-1 enables fine-grained, decoupled control over garment attributes, supporting practical and flexible generation for real-world fashion design applications.
Hyperparameter Analysis. We study the effect of two key hyperparameters in Eq. 1 and Eq. 3: the silhouette guidance weight $\alpha$ and the color conditioning weight $\beta$. As shown in Figure 8, varying $\alpha$ directly impacts the model’s ability to follow the reference silhouette. When $\alpha$ is too low, the generated structure becomes blurry or deviates from the target shape; when too high, it may suppress color and text guidance. We empirically set $\alpha=0.6$ for a balanced structural alignment. Similarly, the color weight $\beta$ controls the influence of the color palette. As $\beta$ increases, color consistency improves steadily, with $\beta=1.0$ yielding visual fidelity performance.
Limitations. While IMAGGarment-1 achieves strong performance in FGG, several limitations remain. First, the base diffusion model’s capability still constrains the visual quality and color fidelity of generated results. Incorporating stronger pre-trained backbones such as Stable Diffusion 3 (Esser et al., 2024) or Flux (Labs, 2024) could further enhance generation quality. Second, the current framework is limited to single-view, flat garment images. However, practical applications in fashion design often require dynamic visualization and 3D consistency. Addressing this requires high-quality multi-view or temporally aligned datasets, which we plan to explore to enable more expressive and realistic garment generation in future work. As Neil Armstrong once said, “That’s one small step for a man, one giant leap for mankind.” We hope this work marks a small but meaningful step toward the future of personalized fashion generation.

### v2

@@ III-C Stage II: Local Enhancement Model
$X=\text{Concat}(C_{g}\otimes C_{m},\ C_{l}),\quad X\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{4}},$ (4)
@@
$x_{t}=\sqrt{\bar{\alpha}_{t}}\cdot x_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (6)
@@
$Z=\text{Concat}(x_{t},\ C_{M},\ X),\quad Z\in\mathbb{R}^{9\times\frac{H}{8}\times\frac{W}{4}}.$ (7)
@@
III-D Training and Inference
Training. The training process is divided into two stages, each targeting a specific set of objectives with separate optimization strategies. We first train the global appearance model independently to generate a semantically coherent garment representation conditioned on silhouette and color. After verifying its performance, we freeze it and train the local enhancement model to inject fine-grained logos guided by spatial masks. This sequential training avoids gradient interference between heterogeneous objectives and ensures each module converges toward its task-specific goal. Both stages adopt mean squared error (MSE) loss to supervise the denoising process.
Stage I. The global appearance model $\theta_{g}$ is trained to synthesize garments that align with the target silhouette and color under textual guidance. To preserve the generative capacity of the pretrained denoising UNet, we freeze all parameters except those of the silhouette UNet and the cross-attention projections in the mixed attention module. Given silhouette features $C_{s}$, text embeddings $C_{t}$, and color features $C_{c}$, we adopt a decoupled training strategy with $L_{\text{silhouette}}$ and $L_{\text{color}}$ losses:
$L_{\text{silhouette}}$ $=\mathbb{E}_{x_{0},\epsilon,C_{t},C_{s},t}\left\|\epsilon-\epsilon_{\theta_{g}}(x_{t},C_{t},C_{s},t)\right\|^{2},$ (8)
@@ III-D Training and Inference
Stage II. The local enhancement model $\theta_{l}$ refines the coarse latent by injecting logos at user-defined locations. To reduce overhead, we fine-tune only the self-attention layers of the logo UNet. Given logo feature $C_{l}$, spatial mask $C_{m}$, and garment latent $C_{g}$, the training objective $L_{\text{logo}}$ is:
$L_{\text{logo}}=\mathbb{E}_{x_{0},\epsilon,C_{l},C_{m},C_{g},t}\left\|\epsilon-\epsilon_{\theta_{l}}(x_{t},C_{l},C_{m},C_{g},t)\right\|^{2},$ (9)
@@
Inference. IMAGGarment supports end-to-end inference through a two-stage pipeline operating in a shared latent space. The global appearance model first generates a latent of coarse garment image conditioned on the input text prompt, silhouette, color, and mask. This process is guided by classifier-free guidance (CFG) [47]:
$\begin{split}\breve{\epsilon}_{\theta_{g}}(x_{t},C_{t},C_{s},C_{c},t)&=w\cdot\epsilon_{\theta_{g}}(x_{t},C_{t},C_{s},C_{c},t)\\ &\quad+(1-w)\cdot\epsilon_{\theta_{g}}(x_{t},t)\end{split}$ (10)
@@
$\begin{split}\breve{\epsilon}_{\theta_{l}}(x_{t},C_{l},C_{m},C_{g},t)&=w\cdot\epsilon_{\theta_{l}}(x_{t},C_{l},C_{m},C_{g},t)\\ &\quad+(1-w)\cdot\epsilon_{\theta_{l}}(x_{t},C_{m},C_{g},t)\end{split}$ (11)
IV Experiments
IV-A Dataset and Metrics
Dataset Construction. As shown in Fig. 4 (a), we construct and release GarmentBench, a large-scale dataset for fine-grained garment generation, containing multi-modal design conditions such as text, sketches, colors, logos, and location masks. It serves as a controllable and extensible benchmark for advancing personalized fashion generation. The construction process is as follows:
(1) Image Collection and Preprocessing. We collect over 189K high-quality garment images from the internet, covering a wide range of categories such as tops, bottoms, and dresses. To eliminate background distractions and focus on the garment region, we apply YOLOv8 [48] for clothing detection and perform tight cropping to obtain clean garment-centric images for further processing.
(2) Text, Sketch, and Color Extraction. For each image, we automatically generate three auxiliary conditions to simulate real-world design guidance: textual descriptions generated by the multi-modal LLM Qwen-VL-Chat [49], covering key attributes such as color, silhouette, and style; structural sketches obtained using Informative-Drawings [50], providing shape and layout priors; and color palettes extracted from single-color garments identified via ResNet50 [51] and clustered using K-means [52].
@@ IV-A Dataset and Metrics
Dataset Description. As shown in Fig. 4 (b), we present representative samples from the GarmentBench dataset, which include fine-grained garment images paired with multi-modal conditions such as textual descriptions, structural silhouettes, color references, logos, and spatial location masks. Additionally, we randomly sample images from the Fashion-ControlNet-Dataset-V3111https://huggingface.co/datasets/Abrumu/Fashion_controlnet_dataset_V3 and apply the same preprocessing pipeline as GarmentBench to construct a test set with 1,267 image-condition pairs for evaluation and comparative analysis.
Dataset Statement. GarmentBench is curated from publicly available fashion imagery under a non-commercial research intent. All personal identifiers were removed; third-party logos and brand marks are included solely to evaluate controllability and remain the property of their respective owners. We release only derived annotations and source URLs (not raw images), together with license notices and a takedown procedure; exact split indices and random seeds are provided for reproducibility.
@@
Evaluation Metrics. We adopt four metrics to comprehensively evaluate visual quality, conditional consistency, and fine-grained controllability. Fréchet inception distance (FID) [53] measures the distribution similarity between generated and real images, reflecting overall realism. Color structure similarity (CSS) [54] assesses the consistency of color distribution, measuring color controllability. Lastly, Logo location accuracy (LLA) [55] quantifies the spatial deviation between generated and target logo positions, reflecting spatial precision. Learned perceptual image patch similarity (LPIPS) [56] reflects human-perceived visual similarity, effectively capturing structural and textural consistency. These metrics comprehensively assess quality and controllability in fine-grained garment generation.
IV-B Implementation Details
@@
In our experiments, both the silhouette UNet and the denoising UNet are initialized with the pretrained Stable Diffusion v1.5 model [^2] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5. The local enhancement model is based on the inpainting variant of Stable Diffusion v1.5333https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-inpainting, with only the self-attention layers being fine-tuned to reduce computational cost. We adopt OpenCLIP ViT-H/14444https://github.com/mlfoundations/open_clip as the CLIP image encoder. All input images are resized to $512\times 640$ resolution. We use the AdamW optimizer [57] with a constant learning rate of $1\times 10^{-5}$. The global appearance model and the local enhancement model are trained for 150K and 50K steps, respectively, using a batch size of 20. During inference, we adopt the DDIM sampler [58] with 50 sampling steps. Unless otherwise specified, the silhouette weight $\alpha$ and color weight $\beta$ in Eq.1 and Eq.3 are set to 0.6 and 1.0. The classifier-free guidance (CFG) scale $w$ in Eq.10 and Eq.11 is set to a default value of 7.0.
@@
TABLE II: Quantitative comparisons on GarmentBench. Ours achieves the top results across all metrics, with best in bold.
Method FID $\downarrow$ CSS $\downarrow$ LLA $\uparrow$ LPIPS $\downarrow$
BLIP-Diffusion* [41] 101.99 104.44 0.13 0.68
ControlNet-Garment* [22] 41.22 83.30 0.36 0.41
AnyDoor* [59] 38.08 68.24 0.65 0.17
IP-Adapter-Garment* [21] 37.95 92.95 0.36 0.43
IMAGGarment (Ours) 17.63 36.16 0.72 0.10
@@
IV-C Baseline Comparisons
Due to the absence of prior work tailored to fine-grained garment generation with multi-condition control, we compare our method against four representative baselines: BLIP-Diffusion [41], AnyDoor [59], ControlNet [22], and IP-Adapter [21]. For subject-driven generation methods, BLIP-Diffusion [41] leverages a learnable Q-Former to align textual and visual embeddings in the latent space, initially designed for subject-preserving generation from text-image pairs. AnyDoor [59] combines identity and detail encoders to reconstruct personalized content, which we adapt to conditions of garment appearance and logo inputs. For plugin-based baselines, we extend ControlNet [22] and IP-Adapter [21] by duplicating and modifying their conditional branches to support multi-conditional inputs, such as silhouette, color, and logo. The adapted versions are referred to as ControlNet-Garment and IP-Adapter-Garment. Specifically, for ControlNet-Garment, we input silhouette, color, logo and mask maps into the ControlNet branch and inject them at each downsampling block, following standard practice. For IP-Adapter-Garment, we extend the official implementation to accept silhouette, color, logo and mask embeddings, which are concatenated and injected via cross-attention. To ensure task relevance, all methods are fine-tuned on our GarmentBench dataset with support for logo-specific conditioning. All methods are trained and evaluated under identical training protocols, input resolutions, and hardware setups. The corresponding quantitative and qualitative results are presented in Table II and Fig. 5, respectively, with detailed analysis provided below.
Figure 5: Qualitative results on seen and unseen GarmentBench samples. The seen set uses original test pairs, while the unseen set involves randomly mixed conditions. IMAGGarment delivers the most consistent outputs, achieving accurate silhouette, color, and logo control across both settings.
TABLE III: Quantitative ablation results on GarmentBench.
Method FID $\downarrow$ CSS $\downarrow$ LLA $\uparrow$ LPIPS $\downarrow$
B0 139.33 104.54 0.15 0.64
B1 47.42 36.65 0.30 0.15
B2 30.19 97.05 0.56 0.33
B3 21.20 43.00 0.65 0.11
B4 46.16 108.25 0.52 0.38
Full 17.63 36.16 0.72 0.10
Figure 6: Qualitative ablation results on GarmentBench. The “Full” configuration achieves the best results, highlighting the importance of each component.
Quantitative Results. As shown in Table II, IMAGGarment achieves the best performance across all four metrics on the GarmentBench dataset, demonstrating its superiority in controllable fine-grained garment generation. Compared to subject-driven methods (BLIP-Diffusion [41], AnyDoor [59]), which rely on global features for personalized reconstruction, IMAGGarment shows substantial improvements in FID, LPIPS, and CSS. These gains highlight the effectiveness of our mixed attention and color adapter modules in achieving coherent multi-condition fusion, resulting in more realistic, perceptually consistent, and color-faithful outputs. In contrast to plugin-based approaches (ControlNet-Garment [22], IP-Adapter-Garment [21]) that simply stack independent conditional branches, IMAGGarment yields significantly higher LLA, reflecting more precise logo placement. Our proposed A3 module drives these improvements, which adaptively injects spatial priors and logo features into the latent space for accurate local control. Overall, these results indicate that global-only conditioning or naive plugin stacking is insufficient for fine-grained control. By contrast, IMAGGarment provides an effective solution for multi-conditional garment synthesis, enabling precise coordination of global structure and local detail.
@@ IV-C Baseline Comparisons
Qualitative Results. Fig. 5 presents qualitative comparisons on both seen and unseen garments. Notably, the seen test set refers to the designated test split of our GarmentBench dataset. In the absence of other suitable public datasets, we assess generalization using an unseen-composition test split constructed by randomly recombining input conditions (e.g., silhouette, color, logo) into combinations that never appear during training, thereby simulating real-world fashion-design scenarios. On seen garments, subject-driven methods (BLIP-Diffusion [41], AnyDoor [59]) reconstruct global appearance but lack spatial control. BLIP-Diffusion retains logo identity yet fails at precise placement due to text-only conditioning, while AnyDoor introduces logo distortions and stylistic artifacts. Plugin-based baselines (ControlNet-Garment [22], IP-Adapter-Garment [21]) treat conditions independently, resulting in poor coordination. ControlNet-Garment suffers from cross-condition interference, and IP-Adapter-Garment often misplaces logos despite preserving structure. In contrast, IMAGGarment achieves accurate control over silhouette, color, and logo placement. On unseen garments, all baselines degrade notably. Subject-driven methods fail to generalize to novel layouts, AnyDoor distorts appearance, and BLIP-Diffusion struggles with logo positioning. Plugin-based methods also falter: ControlNet-Garment produces mismatched outputs, and IP-Adapter-Garment cannot interpret unseen spatial semantics. IMAGGarment remains robust, maintaining alignment across all conditions. This generalization stems from our $A^{3}$ module, which effectively integrates spatial and visual cues in the latent space. These results validate the controllability and flexibility of our method in both seen and unseen settings.
@@
Figure 7: Controllability visualization. Each block varies one input condition while keeping others fixed. Left: Silhouette changes lead to consistent structural adaptation. Middle: Color palette variation results in accurate color transfer. Right: Logo mask adjustment yields precise spatial placement.
IV-D Ablation Study
To validate the effectiveness of each component in our framework, we design a series of ablation variants within the IMAGGarment architecture: B0 uses the vanilla Stable Diffusion v1.5 without any of our proposed modules, serving as the baseline. B1 removes the local enhancement model (Stage II), evaluating the impact of omitting logo injection and spatial control. B2 removes the global appearance model (Stage I), assessing the model’s performance without structured silhouette and color conditioning. B3 removes the color adapter from the global appearance model, isolating the role of color guidance in generation. B4 replaces our mixed attention with vanilla self-attention in the denoising UNet, testing the importance of spatial fusion with silhouette features. Full represents the complete IMAGGarment framework with all proposed modules integrated.
Ablation of Architecture Design. Table III presents the quantitative impact of each component in our proposed IMAGGarment. In B1, which removes the local enhancement stage, the model struggles to place logos precisely, leading to degraded LLA. Although the overall garment structure is preserved, the lack of spatial control prevents accurate logo integration. In B2, without the global appearance stage, the model fails to maintain silhouette and color consistency, resulting in significantly worse FID, LPIPS, and CSS. This demonstrates that local injection alone is insufficient to handle global garment layouts. B3 disables the color adapter, causing notable drops in CSS, highlighting its role in faithful color transfer and control. B4 replaces our mixed attention with standard self-attention, which weakens the fusion of silhouette guidance and causes drops in both LPIPS and FID, indicating reduced realism and structural coherence. The full IMAGGarment achieves the best performance across all metrics, validating the complementary design of each module’s effectiveness in handling multi-condition garment generation. Further, Fig. 6 shows qualitative comparisons. B1 fails to align logos spatially, while B2 produces distorted garments lacking color and silhouette guidance. Despite maintaining logo placement, B3 leads to color mismatch, and B4 generates less coherent garment layouts. In contrast, the full model successfully synthesizes garments with accurate silhouettes, precise logo placement, and faithful color reproduction, demonstrating the benefits of our dual-stage design, color adapter, and mixed attention fusion. Overall, The “Full” configuration achieves the best results, highlighting the importance of each component.